Trying to help my son but it's been waaaay too long ago:
x^2 + y^2 +4x - 6y -3 = 0

Answer by lwsshak3(11628)

x^2 + y^2 +4x - 6y -3 = 0
..
To find out what kind of conic this equation is, we must put it in standard form by completing the square. This is how you do it:
completing the square
(x^2+4x+4)+(y^2-6y+9)=3+4+9=16
(x+2)^2+(y-3)^2=16
This equation is a circle with center at (-2,3) and radius=√16=4
see the graph below as a visual check on the answer
..
y=(16-(x+2)^2)^.5+3